OpenStudy (anonymous):

Estimate the quantity in terms of powers of ten: 2^90

OpenStudy (anonymous):

10^27

OpenStudy (anonymous):

2^90 = 2^(9x10) = (2^9)^10=512^10

OpenStudy (llort):

1.23*10^27

OpenStudy (anonymous):

2^10 is about 10^3 2^20 is about 10^6 2^30 is about 10^9 ... 2^90 is about 10^27